OpenStudy (anonymous):

(1-tanx)^2=sec^2 x-tanx

OpenStudy (anonymous):

I need help verifying it

OpenStudy (anonymous):

the right hand side must be sec^2(x) -2tan(x) , re check, please

OpenStudy (anonymous):

yes i know the answer and I have the steps I just don't understand here I'll show you what I dont understand

OpenStudy (anonymous):

(1+tanx)^2 1-2tanx+tan^2x (1+tan^2x)-2tanx

OpenStudy (anonymous):

yup!! and (1+tan^2x)= sec^2 (x) and it is identity

OpenStudy (anonymous):

yes but how did they get from that 1-2tanx+tan^2 x from (1-tanx)(1+tanx)

OpenStudy (anonymous):

they can't !! but where does it come from? please, don't intermingle the problems.

OpenStudy (anonymous):

* yes but how did they get to 1-2tanx+tan^2 x from (1-tanx)(1+tanx)

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

I don't understand what you ask. The expression is not true 1-2tanx + tan^2 x \(\neq\) (1-tanx) (1+tanx) but 1-2tanx + tan^2 x = (1-tanx) (1-tanx)

OpenStudy (anonymous):

what is the question?? Is the question like " which step is false"

OpenStudy (anonymous):

Your logic is WRONG!!

OpenStudy (anonymous):

no the question is to verify it... I have to lists the steps and I can just go ahead and list them, but I want to understand it

OpenStudy (anonymous):

\[(1+tanx)^2\] \[(1-tanx)(1+tanx)\] I understand up to here THIS STEP IS WRONG!!! you can't go from the first one to the next. if you have \(1~~\color{red}{-}~~tan^2x\) , then it is true. but your expression is (1+tanx)^2

OpenStudy (anonymous):

but where did that 1 come from and how did you get the 2 up there as an exponent

OpenStudy (anonymous):

i got confused yea I thought it went like that (1+tanx)^2 (1-tanx)(1+tanx) but this step i assumed so from here (1+tanx)^2 how do you move on to the next step

OpenStudy (anonymous):

And the original one is not identity, how can you verify. This is counter example. If x = pi/4, then tanx =1 --> left hand side = (1-1)^2 =0 and the right hand side : sec (pi/4)= \(\dfrac{2}{\sqrt2}\) and \(sec^2(pi/4) = 2\) tan (pi/4) =1 therefore the right hand side : \[sec^2(pi/4) - tan (pi/4) = 2-1=1\] so that the left hand side =0 \(\neq\) 1 from the right hand side Conclusion, the expression is NOT true

OpenStudy (anonymous):

yes it is my textbook says it is

OpenStudy (anonymous):

can you scan your text book??

OpenStudy (anonymous):

i screen shotted it its online

OpenStudy (anonymous):

@UnkleRhaukus @ParthKohli @Luigi0210 someone rescues me please

OpenStudy (anonymous):

its 13 it's highlighted pink the first one

OpenStudy (anonymous):

w.t ....???????????? the right hand side is \(sec^2x -\huge 2 \)tanx

OpenStudy (anonymous):

while you post it is \(sec^2 x - tanx\)

OpenStudy (anonymous):

yes this is the answer these are the steps, but I don't understand their processing

OpenStudy (anonymous):

hold on let me check

OpenStudy (anonymous):

sorry i've been doing this all day ... too many numbers

OpenStudy (anonymous):

trash the step to the trashbin!!! just few lines to have the process done (1-tanx)^2 = 1 -2tanx +tan^2 x = sec^2 x -tanx no more!!

OpenStudy (anonymous):

trash what?

OpenStudy (anonymous):

your post is not the good step. I don't care what they do and I don't know why you have to follow their step!! you can prove the problem by your own steps. Period

OpenStudy (anonymous):

(1-tanx)^2 = 1 -2tanx +tan^2 x = sec^2 x -tanx .... that 1 where did you get it from? and that tan^2x ??

OpenStudy (anonymous):

i trashed it

OpenStudy (anonymous):

@lolabieber why do you keep posting the wrong expression?? I told you the right hand side is sec^2 x - 2 tan x, not -tan x

OpenStudy (anonymous):

compare your post and your screenshot

OpenStudy (anonymous):

bc you posted the wrong expression : trash the step to the trashbin!!! just few lines to have the process done (1-tanx)^2 = 1 -2tanx +tan^2 x = sec^2 x -tanx no more!!

OpenStudy (anonymous):

I trash my avatar and replace it by crying baby pic. hahahaha...

OpenStudy (anonymous):

ok, I am sorry for it. , close the post and post it in a new one. OK?? with the screen shot

OpenStudy (anonymous):

yes i Know the -2tanx but can you please tell me what you thought in order to put that 1 -2tanx +tan^2 x ... sorry I'm making you cry imagine me :(

OpenStudy (anonymous):

this post im new here too hah?

OpenStudy (anonymous):

close it.

OpenStudy (anonymous):

in the top right, you see the box "Close" click it

OpenStudy (anonymous):

post another new

OpenStudy (anonymous):

i did and i closed it but it opened again